the perimeter of s rectangle is 58, the length is 5 more then the width what are the dimensions
Answer by TimothyLamb(4379)

2L + 2w = 58
L = w + 5
---
2L + 2w = 58
2(w + 5) + 2w = 58
2w + 10 + 2w = 58
4w = 48
w = 12
L = 17
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---